# BugCraft: End-to-End Crash Bug Reproduction Using LLM Agents in Minecraft

### v1

@@ 4.2.1. Inter-Rater Agreement for Step Synthesizer
Figure 7. Confusion matrix of bug report reproducibility showing inter-rater agreement
4.3. Action Model Analysis
Table 2. Experimental Results and Analysis of Failure Cases
@@ 4.3. Action Model Analysis
Table 2 summarizes the results of our Action Model’s performance on 86 bug reports, conducted as per the methodology in Section 3.3.4.1. Initially, 100 bug reports were considered. However, 14 were deemed ”Irreproducible” during the Step Synthesizer evaluation (see Section 4.2), and an additional 4 were excluded due to errors encountered during the Action Model runs: 3 internal framework crashes and 1 technical issue the potential bug could cause to our setup. These errors prevented the runs from concluding successfully. Therefore, the Action Model ultimately attempted to reproduce 86 bug reports, including those with faulty step clusters identified in the Step Synthesizer analysis. As explained in our evaluation setup (Section 3.3.2), we manually categorized the Action Model’s failures and also specifically inspected the bug reports whose step clusters were generated as ”Faulty”. In these cases, there was a chance for our agent to explore and find the correct way to reproduce the bug by utilizing its inherent knowledge, and the potential feedback it could get from the game.
@@
In our comprehensive end-to-end testing, the Action Model successfully reproduced 26 out of 86 bugs (30.23%). When provided with correct step clusters (”Step Synthesizer Success” cases), the Action Model achieved a 38.60% success rate (22 out of 57). However, when the step clusters were faulty (”Step Synthesizer Failure” cases), the Action Model only managed to recover and reproduce the bug in 4 out of 29 cases (13.79%). These recovery cases are particularly significant because they demonstrate the agent’s ability to overcome errors in the provided plan and successfully reproduce the bug using other available information.
Analyzing the failures when correct step clusters were provided, we found that ”Agent Incapability” was the most common cause (50.88%), with the failures being distributed between ”Poor Decision Making” (29.82%) and ”Stuck in Loop” (19.30%). ”Framework Incapability” accounted for a smaller portion of the failures (7.02%). In cases where the Step Synthesizer failed to generate correct steps, and the Action Model also failed to recover, the vast majority (82.76%) were attributed to the faulty steps.
5. Discussion
This study aimed to evaluate the BugCraft framework through two primary research questions, focusing on the performance of its two main components: the Step Synthesizer and the Action Model. Our findings provide valuable insights into the capabilities and limitations of using LLMs for automated bug reproduction in complex game environments like Minecraft.
5.1. Addressing RQ1
The Step Synthesizer demonstrated a promising ability to transform raw, unstructured bug report data into structured, actionable steps for bug reproduction. It achieved a 66.28% success rate in generating accurate plans (step clusters), as evaluated through the methodology outlined in Section 3.3.1. This indicates that the component is generally effective in understanding the reported issue, extracting relevant information, and synthesizing it into a coherent sequence of actions. The strong correlation between successful plan generation and the Action Model’s success further validates the Step Synthesizer’s performance, highlighting its crucial role in the overall bug reproduction pipeline.
@@ 5.1. Addressing RQ1
However, the 33.72% failure rate reveals that challenges remain. The analysis of faulty plans identified three main error categories: ”Wrong Command” (48.28%), ”Missing Step” (34.49%), and ”Logic Error” (31.03%). ”Wrong Command” errors often stemmed from the LLM relying on outdated command syntax from its pre-training data, especially for newer Minecraft versions. This was particularly evident in the LLM’s struggle to generate correct item tags, which have a detailed schema that changes frequently between versions [^13] https://minecraft.wiki/w/Data_component_format. Additionally, some hallucinated commands were generated that never existed in the game. ”Missing Step” errors indicated that the Step Synthesizer sometimes failed to capture all the crucial details needed for successful reproduction, particularly for complex bugs or those involving recent game mechanics. This often manifested as the LLM failing to adequately track and account for the evolving game state. ”Logic Error” cases demonstrated challenges in ensuring the generated plan’s internal consistency and adherence to game rules.
The Step Synthesizer offers a good foundation for bug reproduction planning, but needs improvement. It requires better integration of version-specific game information, command verification (e.g., syntax parsing), and logical consistency checks. Solutions could include enhancing external data integration and incorporating game environment feedback to refine plans dynamically.
5.2. Addressing RQ2
BugCraft achieved a 30.23% success rate in end-to-end reproduction of crash-inducing bug reports from our curated dataset. This demonstrates a tangible capability for automated game bug reproduction, showcasing the potential of LLMs to bridge the gap between natural language descriptions of bugs and concrete actions within a complex game environment. While this success rate indicates the feasibility of our approach, the corresponding 65.12% failure rate highlights the inherent challenges in automating bug reproduction within such environments.
@@ 5.2. Addressing RQ2
The primary bottleneck in our end-to-end pipeline was the Action Model. Even when provided with accurate step-by-step plans from the Step Synthesizer, the Action Model still failed to reproduce the bug in a majority of cases (56.14%). This underscores the difficulty of translating even well-defined plans into successful in-game actions. Agent incapability was the most significant factor hindering the Action Model’s performance. This category encompassed both ”Poor Decision Making” (29.82%) and being ”Stuck in Loops” (19.30%). ”Poor Decision Making” highlighted the challenges faced by the Action LLM in navigating the complex, visually-rich Minecraft environment. Operating solely on visual input, without a deeper symbolic understanding of the game state, the LLM sometimes struggled to make optimal choices, overlooking environmental cues or deviating from the intended plan. The ”Stuck in Loop” failures further illustrated these challenges, manifesting as ”Menu Loops” (14.04%), ”Command Loops” (3.51%), and even a ”Death Loop” (1.75%). ”Menu Loops,” particularly within the ”Create World” screen, indicated difficulties in reliably navigating complex GUI structures. ”Command Loops” arose from issues in command parsing or execution, with the agent repeatedly attempting unparseable or ineffective commands. The single ”Death Loop” instance was caused by hostile mobs continuously killing the player, preventing progress.
@@
Beyond agent-specific limitations, ”Framework Incapability” contributed to 7.02% of failures when correct steps were available. This suggests that while our custom API provided a foundational level of interaction, it lacked the granularity or robustness needed for certain bug reproduction scenarios. Limitations in precise mouse control, rapid action execution, or interaction with specific UI elements within the game hindered the Action Model’s ability.
Notably, the Action Model successfully reproduced bugs in 13.79% of cases despite faulty plans from the Step Synthesizer. These ”Recovery” cases suggest the agent can leverage its knowledge and context to overcome some inaccuracies, hinting at the potential for a more robust and adaptable Action Model capable of problem-solving. However, the low recovery rate underscores the continued importance of accurate Step Synthesizer plans.
While BugCraft shows promise for automated game bug reproduction, improvements in agent capabilities and framework robustness are essential. Future work should focus on enhancing the agent’s understanding of the game environment, possibly via symbolic or hybrid approaches. Improving the game interaction API for better control/feedback, and adding more robust error detection and recovery, could also boost performance.
6. Threats to Validity
6.1. Construct Validity
@@
A potential threat to construct validity stems from using the Minecraft Wiki for knowledge augmentation. While BugCraft might generalize to other games, especially those with developer support, the exceptionally detailed Minecraft Wiki may not be representative of typical game documentation. This, combined with Minecraft’s popularity, likely led to substantial Minecraft data in the LLM’s pre-training. This pre-existing knowledge could inflate the LLM’s understanding of Minecraft, overestimating BugCraft’s performance. Games with less documentation or popularity might see reduced LLM effectiveness, limiting generalizability. Furthermore, the LLMs might have been trained on the final, corrected versions of Mojira bug reports, even though we used initial, unedited versions, potentially overestimating the LLM’s ability to handle raw bug reports.
6.2. Internal Validity
The internal validity of our study is potentially affected by the inherent stochasticity of LLMs, which typically produce varying outputs even with identical inputs. While we mitigate this by setting the temperature parameter to 0, ensuring deterministic behavior in most components, the Action Model phase presents a unique challenge. During this phase, screenshots of the game are captured, which include Minecraft’s randomly changing, animated background. This unavoidable game feature introduces an element of unpredictability into the Action Model’s processing, as the visual input varies across runs. Further testing is needed to confirm consistency between runs since we were unable to run it multiple times due to resource and time constraints. Furthermore, the limited number of evaluators in the manual analysis phases could introduce bias.
6.3. External Validity
The primary threat to external validity is the limited scope of our evaluation, which focuses solely on crash bugs in Minecraft: Java Edition. This raises concerns about the generalizability of our findings to other types of bugs (e.g., functional bugs, performance issues) and other games. Minecraft, being a sandbox game with a vast and complex environment, may not be representative of all game genres. Furthermore, the specific mechanics and structure of Minecraft: Java Edition, such as the inclusion of an extensive scripting system, might not be comparable to other games. Therefore, while BugCraft demonstrates promising results within the Minecraft context, caution should be exercised when extrapolating these findings to other games or bug types. Further research involving diverse games and bug categories is necessary to establish the broader applicability of our framework. Additionally, our reliance on publicly available bug reports from Mojira might introduce a bias, as these reports may not fully represent the issues encountered by the entire player base.
7. Conclusion
@@
The paper introduces the BugCraft-Bench, a novel benchmark for Minecraft bug reproduction, and BugCraft, an LLM-based framework for end-to-end crash bug reproduction in Minecraft. BugCraft-Bench addresses the dataset gap in game bug reproduction research, offering a curated set of 86 Minecraft crash bug reports for standardized evaluation. BugCraft framework achieves a 30.23% end-to-end success rate in reproducing crash bugs, with its Step Synthesizer demonstrating 66.28% accuracy in plan generation, highlighting LLMs’ potential for automating game bug reproduction. However, limitations in the Action Model, such as the agent getting stuck in loops and poor decision making by the agent were identified, along with Step Synthesizer challenges in command accuracy and detail completeness, especially for newer Minecraft versions.
@@
Moving forward, several key areas warrant further investigation. Future work should focus on enhancing LLM game environment understanding, potentially through collaboration with game developers to create specialized APIs for enhanced in-game movement and interaction, similar to the high-level API Voyager(wang2023voyager, ) uses, called Mineflayer(mineflayer, ). This could also improve bug reproduction across numerous game versions, addressing the limitations of our current vision-based approach. Furthermore, developing comprehensive datasets and robust ground truth verification methods, particularly for non-crash bugs, are crucial. Future benchmarks should incorporate automated or semi-automated ground truth methods to evaluate a wider range of bug types. Assessing BugCraft’s generalizability beyond Minecraft is also essential, requiring investigation into its adaptability to diverse game engines, genres, and complexities through comparative studies. Our roadmap includes improving BugCraft based on identified weaknesses, including addressing challenges with the Action Model and Step Synthesizer. We also plan to extend the agent to handle multiplayer bugs and bugs involving file loading (Worlds, datapacks, world setting JSONs). Ultimately, the advancements presented in this paper provide a strong foundation for future advancements in automated bug reproduction, paving the way for more automatic game development processes.

### v2

@@ IV-B1 Inter-Rater Agreement for Step Synthesizer
Figure 7: Confusion matrix of bug report reproducibility showing inter-rater agreement
IV-C Action Model Analysis
TABLE II: Experimental Results and Analysis of Failure Cases
@@ IV-C Action Model Analysis
Table II summarizes the Action Model’s performance using GPT-4o on 86 bug reports, following the methodology in section III-C4. Initially, 100 reports were considered; however, 14 were deemed ”Irreproducible” (Section IV-B), and 4 Action Model runs encountered errors (3 internal, 1 bug-induced technical issue), leading to 86 attempted reproductions. These included cases with faulty step clusters, where the agent could potentially recover.
@@
In comprehensive end-to-end testing, the Action Model successfully reproduced 26 out of 86 (30.2%) with GPT-4o. When provided with correct step clusters (”Step Synthesizer Success”), the success rate was 38.60% (22 out of 57). Remarkably, when step clusters were faulty (”Step Synthesizer Failure”), the Action Model still recovered and reproduced the bug in 4 out of 29 cases (13.79%), showcasing the agent’s ability to overcome plan errors. Failures with correct plans were primarily due to ”Agent Incapability” (50.88%), encompassing ”Poor Decision Making” (29.82%) and ”Stuck in Loop” (19.30%). ”Framework Incapability” contributed to 7.02% of these failures. In scenarios where the Step Synthesizer provided faulty plans and the Action Model did not recover, the overwhelming majority of failures (82.76%) were attributed to these incorrect steps.
IV-D Comparative Evaluation with Baselines
We evaluated BugCraft against two competitive baselines: OpenAI’s Computer Use Agent (CUA)[33], a recent commercial computer-use model, and UI-TARS-1.5-7B, a strong open-source GUI automation model. All systems were tested on the identical 86 bug reports without any errors from BugCraft-Bench to ensure fair comparison. We use OSWorld’s hyperparameters[28] when running the benchmark.
TABLE III: Baseline Comparison: Performance and Cost Analysis
System Success Active MTTR Cost/
Rate Time Attempt
Human Expert [^13] Active time, the manual effort required to reproduce a bug, measured by the avg. time authors spent during dataset validation. MTTR (Mean Time To Reproduce) is the avg. time for an issue confirmation, calculated from Mojira API. Cost is estimated at \$85/hour (\$64 median wage + 30% employer costs [34]). Success rate is calculated from our percentage-agreement analysis.  83% 20 min 3.41 days \$28.20
BugCraft (GPT-4.1) 34.9% (30) 10.00 min 10.00 min \$1.16
BugCraft (GPT-4o) 30.2% (26) 15.56 min 15.56 min \$1.45
OpenAI CUA 25.5% (22) 6.37 min 6.37 min \$0.65
UI-TARS-1.5-7B 0.0% (0) 3.27 min 3.27 min \$0.02
The experimental results are presented in Table III. BugCraft with GPT-4.1 achieved a 34.9% success rate, outperforming OpenAI CUA (25.5%) and UI-TARS-1.5-7B (0%). The performance improved from 30.2% with GPT-4o to 34.9% with GPT-4.1. Regarding execution efficiency, BugCraft completed reproductions in 10.00 minutes on average with GPT-4.1 (15.56 minutes with GPT-4o), showing a trade-off between success rate and speed compared to CUA’s 6.37 minutes and UI-TARS’s 3.27 minutes. The cost analysis shows that BugCraft operates at \$1.16 per attempt (GPT-4.1), placing it between the more economical CUA (\$0.65) and the low-cost UI-TARS (\$0.02), which reflects the computational demands of its two-stage approach. Notably, while human experts achieve  83% success rate with only 20 minutes of active work, their mean time to reproduce (MTTR) extends to more than this time due to organizational processes, at \$28.20 per attempt based on industry-standard developer costs.
V Discussion
This study evaluated the BugCraft framework via three research questions: step-synthesizer planning (RQ1), end-to-end reproduction (RQ2), and baseline comparison (RQ3).
V-A Addressing RQ1
@@
The Step Synthesizer with GPT-4o demonstrated a promising ability to transform raw, unstructured bug report data into structured, actionable steps for bug reproduction. It achieved a 66.28% success rate in generating accurate plans (step clusters), as evaluated through the methodology outlined in section III-C1. This indicates that the component is generally effective in understanding the reported issue, extracting relevant information, and synthesizing it into a coherent sequence of actions. The strong correlation between successful plan generation and the Action Model’s success further validates the Step Synthesizer’s performance, highlighting its crucial role in the overall bug reproduction pipeline. However, the 33.72% failure rate reveals that challenges remain. The analysis of faulty plans identified three main error categories: ”Wrong Command” (48.28%), ”Missing Step” (34.49%), and ”Logic Error” (31.03%). ”Wrong Command” errors often stemmed from the LLM relying on outdated command syntax from its pre-training data, especially for newer Minecraft versions. This was particularly evident in the LLM’s struggle to generate correct item tags, which have a detailed schema that changes frequently between versions [^14] https://minecraft.wiki/w/Data_component_format. Additionally, some hallucinated commands were generated that never existed in the game. ”Missing Step” errors indicated that the Step Synthesizer sometimes failed to capture all the crucial details needed for successful reproduction, particularly for complex bugs or those involving recent game mechanics. This often manifested as the LLM failing to adequately track and account for the evolving game state. ”Logic Error” cases demonstrated challenges in ensuring the generated plan’s internal consistency and adherence to game rules. The Step Synthesizer offers a good foundation for bug reproduction planning. Integrating version-specific game information, command verification, and logical consistency checks could further improve its reliability. Solutions could include enhancing external data integration and incorporating game environment feedback to refine plans dynamically, however these are challenging to pull off due to their necessity to work in many versions of the game.
V-B Addressing RQ2
@@
BugCraft using GPT-4o achieved a 30.2% success rate in end-to-end reproduction of crash-inducing bug reports from our curated dataset. This demonstrates a tangible capability for automated game bug reproduction, showcasing the potential of LLMs to bridge the gap between natural language descriptions of bugs and concrete actions within a complex game environment. The primary bottleneck in our end-to-end pipeline was the Action Model. Even when provided with accurate step-by-step plans from the Step Synthesizer, the Action Model still failed to reproduce the bug in a majority of cases (56.14%). This is an expected result as even humans can sometimes struggle at reproducing these bugs given a well defined plan. Agent incapability was the most significant factor hindering the Action Model’s performance. This category encompassed both ”Poor Decision Making” (29.82%) and being ”Stuck in Loops” (19.30%). ”Poor Decision Making” highlighted the challenges faced by the Action LLM in navigating the complex, visually-rich Minecraft environment. Operating solely on visual input, without a deeper symbolic understanding of the game state, the LLM sometimes struggled to make optimal choices, overlooking environmental cues or deviating from the intended plan. The ”Stuck in Loop” failures further illustrated these challenges, manifesting as ”Menu Loops” (14.04%), ”Command Loops” (3.51%), and even a ”Death Loop” (1.75%). ”Menu Loops,” particularly within the ”Create World” screen, indicated difficulties in reliably navigating complex GUI structures. ”Command Loops” arose from issues in command parsing or execution, with the agent repeatedly attempting unparseable or ineffective commands. The single ”Death Loop” instance was caused by hostile mobs continuously killing the player, preventing progress.
@@
Beyond agent-specific limitations, ”Framework Incapability” contributed to 7.02% of failures when correct steps were available. This suggests that while our custom API provided a foundational level of interaction, it lacked the granularity or robustness needed for certain bug reproduction scenarios. Limitations in precise mouse control, rapid action execution, or interaction with specific UI elements within the game hindered the Action Model’s ability. Notably, the Action Model successfully reproduced bugs in 13.79% of cases despite faulty plans from the Step Synthesizer. These ”Recovery” cases suggest the agent can explore the given S2R and make concious decisions on the fly to correct the wrong steps. While BugCraft shows promise for automated game bug reproduction, improvements in agent capabilities and framework robustness are essential. Future work should focus on enhancing the agent’s understanding of the game environment, possibly via symbolic or hybrid approaches. Improving the game interaction API for better control/feedback, and adding more robust error detection and recovery, could also boost performance.
V-C Addressing RQ3
V-C1 Isolation of Framework Contributions
Testing against general GUI agents (UI-TARS: 0% success) and specialized computer-use models (OpenAI CUA: 25.5% success) isolates BugCraft’s contributions. The 36.9% success rate, a 32% relative improvement over CUA, stems from three components: (1) Step Synthesizer with game-specific RAG, (2) flexible step clustering, and (3) unified game interaction APIs. These specifically address what generic approaches miss: game domain knowledge and spatial reasoning circumvention.
V-C2 Performance Justification and Efficiency Analysis
While BugCraft’s 34.9% success rate is lower than the 83% achieved by human experts, this overlooks crucial operational advantages. The system’s true value is revealed in its superior efficiency, cost-effectiveness, and scalability. BugCraft demonstrates remarkable time efficiency, completing a reproduction in 10 minutes compared to a human’s 3.41 day turnaround, which is often delayed by organizational overhead. This vast reduction in time allows our system to be suitable for rapid triage. Financially, the system offers a 10-fold cost reduction per successful reproduction-\$3.44 for BugCraft versus \$35.25 for a human - and its success with crash bugs is definitively verifiable. Finally, unlike human experts who work sequentially, the framework processes bugs in parallel across multiple instances. This capability is critical for projects like Minecraft that handle hundreds of bug reports weekly, allowing for a high-throughput, automated workflow that humans cannot match.
V-C3 Technical Differentiation from Prior Work
BugCraft differs from prior mobile/web GUI automation through three game-specific innovations: (1) intelligent bug analysis by Step Synthesizer, (2) RAG integration with game knowledge bases to disambiguate vague reports, and (3) unified macro API design handling both 2D menus and 3D gameplay. These address challenges unique to games, spatial reasoning, complex state dependencies, and version-specific mechanics, that 2D GUI frameworks cannot handle.
Figure 8: Case study analysis of MC-161902 comparing how our framework successfully places the water block whereas OpenAI-CUA fails.
Case Study: We analyzed bug MC-161902151515https://bugs.mojang.com/browse/MC-161902, a game crash triggered when specific mobs contact water, to highlight a critical difference between BugCraft and baseline models. The core challenge to reproduce this bug is placing the water as shown in Figure 8.
  1. 1.
    OpenAI CUA: Despite being encouraged to use commands, the model opts for a physical approach. It successfully gets a water bucket but fails when trying to aim and place the water on a valid surface, as it is unable to see the difference and fails to reproduce the bug.
  2. 2.
    BugCraft: Our Step Synthesizer analyzes the bug report and identifies the most efficient path to the goal. It intelligently bypasses the difficult physical task by using a command to spawn the water block directly. The Action Model then easily follows the steps and is able to reproduce the bug.
V-C4 Qualitative Analysis of the Models
OpenAI-CUA navigates menus better than our Action Model, but its poorer solution choices reveal the value of our Step Synthesizer’s bug analysis. UI-Tars-1.5-7B fails outright; as a very small model specialized for computer GUI usage, it does not know Minecraft well and gets lost in the menus or idles in-game. The authors also point out the fact that the smaller model is not optimized for game specific scenarios [^16] https://github.com/bytedance/UI-TARS#limitations. GPT-4.1 achieved a higher end-to-end success rate than GPT-4o (34.9% vs. 30.2%; 30 vs. 26 of 86). A qualitative execution-log analysis indicates GPT-4.1 followed the prescribed step clusters more closely, whereas GPT-4o exhibited more off-plan attempts. The models’ successes only partially overlapped (20 both; 10 only GPT-4.1; 6 only GPT-4o; McNemar p=0.45), yielding an oracle coverage of 36/86 (41.9%). This complementarity implies that mixing agent strategies can increase coverage in bug reproduction.
VI Threats to Validity
VI-A Construct Validity
@@
A potential threat to construct validity stems from using the Minecraft Wiki for knowledge augmentation. While BugCraft might generalize to other games, especially those with developer support, the exceptionally detailed Minecraft Wiki may not be representative of typical game documentation. This, combined with Minecraft’s popularity, likely led to substantial Minecraft data in the LLM’s pre-training. Games with less documentation or popularity might see reduced LLM effectiveness. To address this, we allow the Step Synthesizer to be augmented with any kind of external knowledge. To directly address the concern that LLMs might have been trained on final, corrected versions of Mojira bug reports, we made a conscious decision to use only the initial, unedited versions. We used the smaller UI-Tars-1.5-7B model, as the larger one was not publicly available. Despite following the official implementation, since small models are sensitive to prompt changes, our results may be suboptimal as we lacked access to potentially better internal prompt templates.
VI-B Internal Validity
For internal validity, we considered the inherent stochasticity of LLMs. To promote consistency and deterministic behavior in most components, we mitigated this by setting the temperature parameter to zero. While this is an unavoidable aspect of the game environment and a challenge for vision-based agents, we acknowledge that resource and time constraints limited our ability to conduct multiple runs for every bug report to fully quantify its impact in this study. Furthermore, the limited number of evaluators in the manual analysis phases could introduce bias.
VI-C External Validity
A key consideration for external validity is that our evaluation is limited to crash bugs within Minecraft: Java Edition, which raises questions about direct generalizability to other bug types or different games. However, it is worth noting that the size of our evaluation dataset is comparable to automated bug reproduction frameworks in other domains-for instance, similar Android bug reproduction frameworks have been evaluated on datasets sized similarly to ours [14, 15]. Moreover, BugCraft demonstrated its adaptability by successfully processing reports from over 70 different Minecraft versions, spanning more than six years of the game’s development. Minecraft itself, with its vast and evolving environment, provides a robust test case for our approach due to its inherent complexity. While these results within Minecraft are promising, caution is still warranted when extrapolating directly to all other bug types or game genres, highlighting the need for future research to explore BugCraft’s performance in diverse contexts. Additionally, our reliance on publicly available Mojira bug reports, while transparent, might not fully represent the entire spectrum of issues encountered by all players.
VII Conclusion
@@
The paper introduces the BugCraft-Bench, a novel benchmark for Minecraft bug reproduction, and BugCraft, an LLM-based framework for end-to-end crash bug reproduction in Minecraft. BugCraft-Bench addresses the dataset gap in game bug reproduction research, offering a curated set of 86 Minecraft crash bug reports for standardized evaluation. BugCraft framework achieves a 34.9% end-to-end success rate in reproducing crash bugs, outperforming existing computer-use approaches, highlighting LLMs’ potential for automating game bug reproduction. However, limitations in the Action Model, such as the agent getting stuck in loops and poor decision making by the agent were identified, along with Step Synthesizer challenges in command accuracy and detail completeness, especially for newer Minecraft versions.
@@
Moving forward, several key areas warrant further investigation. Future work should focus on enhancing LLM game environment understanding, potentially through collaboration with game developers to create specialized APIs for enhanced in-game movement and interaction, similar to the high-level API Voyager [23] uses, called Mineflayer [35]. This could also improve bug reproduction across numerous game versions, addressing the limitations of our current vision-only based approach. Developing comprehensive datasets and robust ground truth verification methods, especially for non-crash bugs, is crucial, requiring future benchmarks to incorporate automated or semi-automated verification tools. Finally, assessing BugCraft’s generalizability is essential, necessitating comparative studies across diverse game engines, genres, and complexities. Our roadmap focuses on improving BugCraft’s Action Model and Step Synthesizer based on identified weaknesses, plus adding support for multiplayer and file-loading bugs (worlds, datapacks, JSON settings). Ultimately, the advancements presented in this paper provide a strong foundation for future advancements in automated bug reproduction, paving the way for more automatic game development processes.